How much heat is required to change 200ml of ice at - 22°C(at typical freezer temperature) into steam? (10 marks)


1
Expert's answer
2021-04-13T06:28:19-0400

Answer

the amount of heat required to convert ice at 0° to water at 0° This can be calculated as

"Q_1= 334 kJ/kg\times200\times0.001kg\\=66.8kJ"

where Latent heat of fusion of water =334kJ/kg

Now the amount of heat required to raise the temperature of water from zero degree centigrade to 20 degree centigrade.

"Q_2= 2kg * 4.186 kJ/kg °C\times(-22°C)" = -184.18kJ

whereSpecific heat of water = 4.186 kJ/kg °C

Hence total heat required

"= Q_1 + Q_2\\ = 334-184.18\\=149.81kJ"